An A-frame cabin has a triangular front with a base of 30 ft and a
height of 20 ft. If the front is to be glass that costs $3 per square foot, what will the
glass cost?

Answer by checkley77(12844)

A=BH/2
A=30*20/2=600/2=300 FT^2.
300*3=$900 COST OF THE GLASS.